OpenStudy (anonymous):

Solve for x: 5|2x - 2| + 8 = 18 x = 2, x = -2 x = -2, x = 0 x = 2, x = 0 x = 0, x = 1 over 2

OpenStudy (campbell_st):

start by subtracting 8 from both sides of the equation and you get 5|2x -2| = 10 divide both sides by 5 |2x - 2| = 2 now you need to solve 2 equations 2x - 2 = 2 and 2x - 2 = -2 hope this helps

OpenStudy (anonymous):

Is it A because the first one is 2

OpenStudy (anonymous):

and the seconds one is negative